Two blocks A and B are connected by a light flexible chord which passes over a frictionless pulley. The mass of block is 3kg. Block A rests on an inclined plane whose angle of inclination is 35 degrees. What must be the mass of block B so that the system will be in equilibrium? There is no friction anywhere in the system



1
Expert's answer
2022-02-04T07:41:46-0500

The net force must be zero for the blocks to be in equilibrium:


"T-m_Bg=T-m_Ag\sin\theta,\\\space\\
m_A=\dfrac{m_B}{\sin\theta}=5.23\text{ kg}."